Triangles ABC and DEF are similar. Find the measures of the remaining sides a=8, c=10, b=6, f=9
Answer by venugopalramana(3286)

Triangles ABC and DEF are similar. Find the measures of the remaining sides a=8, c=10, b=6, f=9
USING STANDARD NOTATION WE HAVE
a/d=b/e=c/f
8/d=6/e=10/9
8/d=10/9
10d=8*9=72
d=72/10=7.2...similarly
10e=54
e=5.4